Person A takes 6 days less than the time taken by the person B to finish a piece of work. If both A and B together can finish it in 4 days. Find the time that B would take to finish this work by himself.
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Person A takes 6 days less than the time taken by the person B to finish a piece of work. If both A and B together can finish it in 4 days. Find the time that B would take to finish this work by himself.
------------------
B DATA:
time = x day/job ; rate = 1/x job/day
---
A DATA:
time = x-6 days/job ; rate = 1/(x-6) job/day
------
Together DATA:
time = 4 days/job ; rate = 1/4 job/day
-------
Equation:
rate + rate = together rate
1/x + 1/(x-6) = 1/4
------
4(x-6) + 4x = x(x-6)
8x-24 = x^2 - 6x
x^2 - 14x + 24 = 0
------
x = [14 +-sqrt(14^2 - 4*1*24)]/2
x = 12 days (time for "A" to do the job)
x-6 = 6 days (time for B to do the job alone)

Answer by ikleyn(52784)

.
To see many other similar solved problems, look into the lesson
    - Using quadratic equations to solve word problems on joint work
in this site.

-----------------
There is a bunch of lessons on Joint work problems
    - Rate of work problems
    - Using Fractions to solve word problems on joint work
    - Solving more complicated word problems on joint work
    - Using quadratic equations to solve word problems on joint work
    - Solving rate of work problem by reducing to a system of linear equations
    - Selected joint-work word problems from the archive
    - Joint-work problems for 3 participants
    - Had there were more workers, the job would be completed sooner
    - One unusual joint work problem
    - Special joint work problems that admit and require an alternative solution method
    - Joint work word problem for the day of April, 1
    - OVERVIEW of lessons on rate-of-work problems
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic
"Rate of work and joint work problems"  of the section  "Word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.